Prev Next

SW&IT Meeting Minutes 2022-02-07

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220124
LR20220124 - Bec, Matthieu 80.0
LR20220124 - Mechtley, Brandon 40.0
LR20220124 - Eychaner, Glenn 80.0
LR20220124 - Edmeades, David 80.0
LR20220124 - Grenz, Paul 20.0
LR20220124 - Hooper, Stephen 0.0
LR20220124 - Zhang, Yang 80.0
LR20220124 - Zhang, Xianyu 85.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management, meet Grant West
    • DMS archive index, mongo refresher
    • shark-nir doc review
    • LN prep and status meeting
    • PR loadrecmat padding fix
  • This week
    • ATT PM
    • archive meeting
    • set_z0 follow-up
    • all-hands presentation Feb-18
    • DMS archive mongo, lit2/webpack
    • PB work preps readiness

YZ

  • Previous:
    • Updated lib-telemetry (v4.2) on TCS machines (seems to be working okay).
    • Several MODS issues (ongoing):
      • Partner accounts passwords on MODS machines (or are passwords disabled?).
      • Rework certain MODS utilities on Fedora workstations to distinguish partner accounts.
      • Some MODS WCS/archive requirements discussion with Olga.
      • Port "dispMODSx2d" to Fedora workstation.
      • Review MODS tcs client.
      • Working on cleaning additional data on MODS data machines.
    • TMS meeting and software work.
    • Safety training.
  • Next:
    • New GCS patch requested by John (to help with his IT#8315 investigation).
    • Possible TMS daytime tests.
    • Possible science archive meeting.
    • Continue to work on various MODS issues outlined earlier.
    • TMS meeting and continued software work.
    • Finish remaining safety training.

DE

  • Vault swtichover
    • Ultimately, it went fine. I'm not pleased that my prep was based on an incorrect understanding of the workings, but the actual system is functional
    • Key users have installed Client 2020, no complaints. In fact, JH reports that 2020 has solved an annoyance he had with 2018 on his laptop.
    • PM started with Ketiv. Jose kicked back at me, claiming that he'd have taken care of the license server, etc. I'm still not cool with not having had all of the info beforehand.
      • 2018 was never going to work past 1 Feb. I don't know why I was allowed to think otherwise
      • Was this just a style mismatch? Do most of their clients go totally hands-off?
      • Lessons learned for dealing with them in the future. I will make sure we have a complete understanding of what is needed and who will do what.
    • Now that we're moved over to the new VM, new Server software, etc. and I understand the licensing I can license clients up to 2022. We need to refresh the liceneses every year now, so I'll experiment with how that works and set up a calendar reminder with enough leeway to make sure that my edu credentials are also in good standing.
  • 507 Polycom ticket
    • I checked that night, and everything looked like it was working fine. Mountain Polycom was on, clients were on. I haven't heard from Alex about trying again with the instructions to connect to the Zoom bridge.
    • I thought I'd left instructions in there, but that was long ago and might have been lost.
  • VPN requests
    • Carmelo: As I said this morning to AC, our VPN works and I was even able to connect from my computer with his credentials. Something is probably going on with his local OS config.
    • Delete old user, add new one for JC@ND.
  • Next
    • Veeam licensing poke. Again.
    • VMWare licensing check
    • Backup health check after VM volume was close to full
    • Work out SMB connection from new Vault VM to new Synology
      • Double extra check other shares re: security implications of enabling guest for Vault backup

GE

  • Last week
    • Fixed a few more bugs in the RTD discovered during testing.
    • Found that 'ssh -YC' is worse than X2Go for running LUCI at Steward or on the mountain.
    • Merged the new RTD, and deployed to luci.luci and the observer computers for further testing by SciOps.
    • Began work on the observer GUI; IT 6138 (problem with pauses in GUI) and IT 8529 (GUI possibly disconnecting from luci.luci).
  • This week
    • Continue testing the new RTD; see the notes in the pull requests for details.
    • Continue work on the observer GUI; IT 6138 (problem with pauses in GUI) and IT 8529 (GUI possibly disconnecting from luci.luci).
    • Examine the other LUCI GUIs to check for GUI modifications from threads other than the Event Dispatch Thread.

BM

  • Last week:
    • Merged in Alfio's gain optimization changes (use max_iterations), installed on soul-sx
    • Finished setting up ao-supervisor on dx machines (deploy keys etc.)
    • ICE dispatch debugging.
    • AOG meeting, getting caught up on set_z0 logic/use of banks.
    • Updated config-gain-ramp to ignore gain ramp if delay <= 0
    • Some work with ao-confcalib git.
  • This week:
    • merge in LoadRecMat padding fix
    • ICE dispatch bug
    • PB prep
    • so-confcalib git
    • SOUL S/W checkout/testing (skipframe button)
    • LBTI S/W checkout/testing

SH

  • Moved archive and repository to new Synology
    • telemetry in progress as of this morning
  • AO switch work
  • Room 530 monitor work
  • Safety Committee
  • Mike L. network stuff
  • scienceops web page hosting

next:

  • Moving stuff to new Synology
  • Upgrading rm507 computer to Fedora 34
  • VM2 in Tucson
    • VM licensing

PSG

  • Last week
    • Testing reset.
    • Reset worked, saw errors and was able to clear them.
  • This week
    • Checking in fixes for GCS
Topic revision: r9 - 08 Feb 2022, PaulGrenz
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback